OpenStudy (anonymous):

I don't want a direct answer, I want someone to help me understand and walk me through the question. Question: Bo factored x^2 – 10xy + 16y^2 as follows: Line 1 x2 – 10xy + 16y2 Line 2 x2 – 8xy – 2xy + 16y2 Line 3 x(x – 8) – 2y(x – 8y) Line 4 prime A. Did Bo correctly factor the polynomial? Justify your response by explaining why it is is factored correctly or by showing the correct solution if it is not factored correctly. B. Explain how you would know if a polynomial of the form ax^2 + bxy + cy^2 was prime. Someone help plz, I'll give medals.

OpenStudy (anonymous):

@triciaal @zepdrix @mathmale @mathstudent55 @SolomonZelman

OpenStudy (tygrr321):

idk srry

OpenStudy (anonymous):

@tygrr321 Alright, that's fine. Thanks anyways.

OpenStudy (anonymous):

@ParthKohli @imqwerty @Directrix @robtobey @SolomonZelman

Directrix (directrix):

When x2 – 8xy was factored, the Bo got x ( x - 8). What happened to the y in the term -8xy ? What should Bo get?

OpenStudy (anonymous):

@Directrix Bo should have gotten x(x - 8y)?

Directrix (directrix):

x2 – 8xy = x ( x - 8y) That is correct. @Needhelpstudying

OpenStudy (anonymous):

Hooray!

OpenStudy (aihberkhan):

a) Okay... the correct way is \[x(x – 8y) – 2y(x – 8y)\] Due to this bo did not factor correctly. So, the best solution would be: \[x^2 – 10xy + 16y^2 = (x - 2y) (x - 8y)\] b) If both a and c are prime then the polynomial is prime. For example, \[X^2+5\] is prime. Do you understand? But, overall I think @Directrix will help completely. I just wanted to add in my thoughts :) But, @Directrix will explain better :)

OpenStudy (aihberkhan):

Hope this helped! Have a great day and Happy Holidays! :) Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @Needhelpstudying

Directrix (directrix):

@Needhelpstudying Start at line 2 and work the problem for Bo. Line 2 x2 – 8xy – 2xy + 16y2

OpenStudy (anonymous):

@Directrix Alright

OpenStudy (anonymous):

x2 - 8xy - 2xy + 16y2 = (x2 - 8xy) - (2xy + 16y2)

OpenStudy (anonymous):

Find GCF: GCF for first binomial: x GCF for second binomial: 2y Factor out GCF: x(x - 8y) - 2y(x - 8y)

OpenStudy (anonymous):

Combine binomials: (x - 2y)(x - 8y) = FINAL Answer

Directrix (directrix):

Line 2 x2 – 8xy – 2xy + 16y2 Line 3: x( x - 8y) - 2 y ( x - 8y) (x - 8y) is a common factor Line 4: ( x - 8y) ( x - 2y)

Directrix (directrix):

Correct.

Directrix (directrix):

What about this: B. Explain how you would know if a polynomial of the form ax^2 + bxy + cy^2 was prime.

OpenStudy (anonymous):

Alright, so now I know that Bo didn't factor the polynomial correctly, and I know what the correct answer is. Part A is done

OpenStudy (anonymous):

@Directrix What exactly does a prime polynomial mean?

Directrix (directrix):

Yes and part B is up to solve.

Directrix (directrix):

A prime number such as 3 will not factor (other than 1*3). A prime polynomial will not factor.

OpenStudy (anonymous):

@Directrix Sorry, my laptop shut off

OpenStudy (anonymous):

@Directrix B. Explain how you would know if a polynomial of the form ax^2 + bxy + cy^2 was prime. I would know if a polynomial of the form ax^2 + bxy + cy^2 was prime, because it would not factor at all/any further. Would that answer be correct?

OpenStudy (anonymous):

@Directrix Hello??

OpenStudy (anonymous):

@Directrix Was I correct?

Directrix (directrix):

>I would know if a polynomial of the form ax^2 + bxy + cy^2 was prime, because it would not factor at all/any further. While that is true, I think that the question is asking for a more detailed answer.

OpenStudy (anonymous):

@Directrix Like give an example of a prime polynomial and show that it's prime?

Directrix (directrix):

@AihberKhan posted this above: >>b) If both a and c are prime then the polynomial is prime. I saw that over at Yahoo answers but there was no explanation so I do not know if it is true.

OpenStudy (anonymous):

@Directrix Any way you can find out if it's true?

Directrix (directrix):

We could test it by seeing what happens when factoring a polynomial where a and c are prime. 3x^2 + 5x - 2 3 is prime and 2 is prime. Can you factor: 3x^2 + 5x - 2 @Needhelpstudying

OpenStudy (anonymous):

Yea sure

OpenStudy (anonymous):

3x^2 + 5x - 2 = 3x^2 + 3x + 2x - 2 Group them together: (3x^2 + 3x) + (2x - 2) Find GCF GCF of first binomial is 3x GCF of second binomial is 2 Divide GCF 3x(x + 1) + 2(x - 1)

OpenStudy (anonymous):

Add terms together: (3x + 2)(x + 1)(x - 1) Wait... wut?

OpenStudy (anonymous):

@Directrix help

Directrix (directrix):

Did an online calculator do the factoring you posted? :)

OpenStudy (anonymous):

No, but I did have to use a calculator to find out what -2 / 2 was. (My brain turned off)

Directrix (directrix):

3x^2 + 5x - 2 Find two numbers that multiply to a*c and add to b a = 3 c = -2 b = 5 Find two numbers that multiply to -6 and add to 5 @Needhelpstudying

OpenStudy (anonymous):

Why?

Directrix (directrix):

>>Why? That is the procedure for factoring by grouping.' 3x^2 + 5x - 2 = (x+2)(3x-1) 3 and 2 are prime so I don't know what to say about @AihberKhan 's post that >>"b) If both a and c are prime then the polynomial is prime." It seems not to be true in this case.

OpenStudy (anonymous):

Do you think my teacher will mind if all I say is something like "I would know if a polynomial of the form ax^2 + bxy + cy^2 was prime, because it would not factor at all/any further. "

OpenStudy (anonymous):

@Directrix Cause technically thats what theyre asking

Directrix (directrix):

I see your point. I don't know for sure what the teacher will accept. Go with it and see.

OpenStudy (anonymous):

Alright. Thanks for all your help on this @Directrix !!!!

Directrix (directrix):

@AihberKhan Would you explain this. Thanks. > b) If both a and c are prime then the polynomial is prime.

OpenStudy (anonymous):

brb

OpenStudy (aihberkhan):

Okay. Let me explain part b "_ So basically if a in the equation and c in the equation are prime... this simply means that the whole equation is prime.. As I said in my answer... for example... \[x^2 + 5\] is prime. So if a and c look like this it is prime.. You can then conclude that the entire equation is prime.. Hope this helped further explain! :) Sorry of this doesn't helped I don't know how I can explain it any further but good luck anyways!! @Needhelpstudying Have a great day and Happy Holidays! :)

Directrix (directrix):

@AihberKhan What about this where a and c are prime? 3x^2 + 5x - 2 = (x+2)(3x-1) 3 and 2 are prime so I don't know what to say about @AihberKhan 's post that >>"b) If both a and c are prime then the polynomial is prime." It seems not to be true in this case.

OpenStudy (anonymous):

@Directrix While you guys sort this out, ima write my answer for part a.

OpenStudy (aihberkhan):

Oh I see. I am truly sorry, I did not mean to give false information by any means! Thank you so much @Directrix for correcting me! I truly apologize for giving any false information... just wanted to add in a little thought. Thank you @Directrix for correcting me! I will be sure to think about that in the future! :)

imqwerty (imqwerty):

:) ima give the proof why the polynomial will be considered prime is \(a\) and \(c\) were prime lets say the polynomial is- \(ac^2+bx+c\) we know that the product of roots is given by->\(\large\frac{c}{a}\) lets say that the roots are-\(\alpha\) and \(\beta\) if both \(c\) and \(a\) are prime then - \(\large \alpha \times \beta =\frac{c}{a}\) now because a and c are prime either of \(\alpha\) or \(\beta\) or both will have to be in fraction form and we know that the polynomial \(ax^2+bx+c\) can be when written in factored form is this- \((x-\alpha)(x-\beta)\) since the roots are in fraction form the factors are not integers and thus the polynomial has no integer divisors so it will be a prime polynomial

Directrix (directrix):

No need to apologize. @AihberKhan We are all just thinking about this problem.

OpenStudy (aihberkhan):

Oh okay great! :) @imqwerty

imqwerty (imqwerty):

:) okay now for part A he did the mistake in 2nd step he used this method-> splitting the middle term and he did it wrong reason why he did it wrong-> lets say the polynomial is -> \(ax^2+bx+c\) when we split the middle term \(b\) we split it in such a way that the product of the 2 things into which we split our middle term must be equal to this product- \(ax^2 \times c\) and the sum of the two things into which we split must be equal to \(bx\)

OpenStudy (anonymous):

Alright guys, thanks for all your help on this

imqwerty (imqwerty):

for part B-> the polynomial is prime when the factors are not integral :)

imqwerty (imqwerty):

and we know what is the situation when this occurs :) when a and c both become prime